OpenStudy (anonymous):

x=2y-7 4x+5y=24 A.(7,3) B.(3,7) C.(1,4) D.(4,1)

OpenStudy (anonymous):

@DarkMoonZ

OpenStudy (mvanessa):

It's C Because... 4(2y - 7) + 5y = 24 8y - 28 + 5y = 24 13y - 28 = 24 13y = 52 y = 4 Finally... x = 2(4) - 7 = 1 Solution: (1,4) I hope this helps!